2 is to 55 as 6 is to ?
Answer by stanbon(75887)

2 is to 55 as 6 is to ?
---------
6/x = 2/55
---
Invert::
x/6 = 55/2
----
x = 330/2
----
x = 165
-------------
Cheers,
Stan H.